OpenStudy (anonymous):

Functions f and g are defined by f(x): k-x for xER, where k is a constant. g(x): 9/(x+2) for xER, x does not = -2 Find the values of k for which the equation f(x) = g(x) has two equal roots and solve the equation f(x) = g(x) in these cases.

OpenStudy (anonymous):

k-x=9/(x+2) (k-x)(x+2)=9 kx+2k-x^2-2x=9 -x^2+(k-2)x+2k-9=0 ----a a=-1 b=(k-2) c=(2k-9) b^2-4ac=0 (k-2)^2-4(-1)(2k-9)=0 solve the above equation and find k put it back in a and find x

OpenStudy (anonymous):

Thanks! Very much appreciated.

OpenStudy (anonymous):

np :)